Problem Set 1A/1: Trade Do this problem set if your surmame (family name) begine R-Z Name: The small nations of Arboc and Arbez can produce two goods, wheat and coffee. The maximum quantity of each good is shown by the production possibility information below. Assume the two nations trade in a world containing many small nations. Costs are constant within each country. Arboc Arbez wheat 50 0 wheat 30 0 0 0 90 an and and and 100 1. What is the opportunity cost of one unit of wheat in Arboc? 2. 3. 4. 5. 6. What is the opportunity cost of one unit of wheat in Arbez? Which country has a comparative advantage in the production of wheat? What is the opportunity cost of one unit of coffee in Arboc? What is the opportunity cost of one unit of coffee in Arbez? Which country has a comparative advantage in the production of coffee? The terms of trade are 1 unit of coffee per unit of wheat. Arboc should produce which good? 7. 8. The terms of trade are 2.5 units of coffee per unit of wheat. Arboc should produce which good? Arbez should produce which good? 9. The terms of trade are 4 units of coffee per unit of wheat. Arbez should produce which good? 10. The terms of trade are 3 units of coffee per unit of wheat. Which country could not gain from trade?

Problems 1-10?
Answer #1

1. Opportunity cost of one unit of wheat in Arboc = Total coffee sacrificed in Arboc/Total wheat produced in Arboc = 100/50 = 2 units of coffee

2. Opportunity cost of one unit of wheat in Arbez = Total coffee sacrificed in Arbez/Total wheat produced in Arbez = 90/30 = 3 units of coffee

3. As opportunity cost of one unit of wheat in Arboc is less than opportunity cost of one unit of wheat in Arbez, so Arboc has a comparative advantage in the production of wheat.

4. Opportunity cost of one unit of coffee in Arboc = Total wheat sacrificed in Arboc/Total coffee produced in Arboc = 50/100 = 0.5 units of wheat

5. Opportunity cost of one unit of coffee in Arbez = Total wheat sacrificed in Arbez/Total coffee produced in Arbez = 30/90 = 0.33 units of wheat

6. As opportunity cost of one unit of coffee in Arbez is less than opportunity cost of one unit of coffee in Arboc, so Arbez has a comparative advantage in the production of coffee.

Note: Post four subparts at a time.